OpenStudy (anonymous):

Write an equation for a circle with center at (2,3) and radius of 6

OpenStudy (sepeario):

(x - 0)² + (y - 0)² = 6² x² + y² = 36

OpenStudy (anonymous):

alrighty so it is x^2+y^2=36 :) thanks

OpenStudy (aakashsudhakar):

(h, k) --> (2, 3) r = 6 The equation for a circle is... (x - h)^2 + (y - k)^2 = r^2 Plug in your values to get your circle equation. (x - h)^2 + (y - k)^2 = r^2 (x - 2)^2 + (y - 3)^2 = 36 ^^ That's a satisfactory answer, or you can even get fancy if you like. (x - 2)^2 + (y - 3)^2 = 36 (x^2 - 4x + 4) + (y^2 - 6x + 9) = 36 x^2 - 4x + y^2 - 6x + 13 = 36 x^2 - 4x + y^2 - 6x = 23 ^^ That's also acceptable.

OpenStudy (aakashsudhakar):

x^2 + y^2 = 36 is not the correct equation for the circle.

OpenStudy (anonymous):

(x-2)^2+(y-3)^2=36

OpenStudy (aakashsudhakar):

Yes, that's correct.

OpenStudy (anonymous):

alright thanks